OpenStudy (anonymous):

Which expression is equivalent to 34 ⋅ 3−9? 1 over 3 to the 13th power 1 over 3 to the 5th power 35 313

OpenStudy (anonymous):

@Michele_Laino @Nnesha

OpenStudy (anonymous):

3^4 x 3^-9

OpenStudy (anonymous):

c and d are 3^5 and 3^13

OpenStudy (ribhu):

3^-5

OpenStudy (michele_laino):

hint: \[\Large {3^4} \cdot {3^{ - 9}} = {3^{4 - 9}} = ...?\]

OpenStudy (anonymous):

3^-5?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

how is that the answer?

OpenStudy (michele_laino):

we can write this: \[\Large {3^4} \cdot {3^{ - 9}} = {3^{4 - 9}} = {3^{ - 5}} = \frac{1}{{{3^5}}}\]

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

ok thanks i still have some questions

OpenStudy (michele_laino):

ok!